# Quantum Magnetic Skyrmions on Kondo-type Lattices

### v3

@@ Abstract
Next, to discuss the role of the i-electrons, we move to a spin-electron (sp-el) and a spin-only (sp) system in a rhombus-shaped $11\times 1$ 11 finite triangular lattice, also treated with OBC. We again compare their respective ground states using $\eta^{\prime}=N^{-1}\sum_{i\in C}\big{|}\langle\hat{\mathbf{S}}_{i}^{\text{sp-el}}\rangle-\langle\hat{\mathbf{S}}_{i}^{\text{sp}}\rangle\big{|}\big{/}\big{|}\langle\hat{\mathbf{S}}_{i}^{\text{sp}}\rangle\big{|}$. We choose $D\!\!=\!\!0.2,J\!\!=\!\!0.15,B^{z}\!\!=\!\!0.06,t_{h}\!\!=\!\!1,g\!\!=\!\!2$, for which both systems exhibit a single-skyrmion texture. This is confirmed by the value of the quantum scalar chirality [42] $Q=\pi^{-1}\sum_{<ijk>\in C}\langle\hat{\mathbf{S}}_{i}\cdot[\hat{\mathbf{S}}_{j}\times\hat{\mathbf{S}}_{k}]\rangle$, where the sum runs over all non overlapping triangles formed by neighboring sites $i$, $j$, $k$. For $S=1/2$ l-spins, $|Q|\sim 1$ signals the presence of a skyrmion. For the case at hand we respectively find $Q_{\text{sp}}^{\text{sp-el}}=-0.618$, and $Q^{\text{sp}}=-0.624$. Even in this strong Kondo regime ($g/t_{h}=2$), the mean-field electron-spin expectation values $\langle\hat{\mathbf{s}}_{i}\rangle$ are only few percents of the corresponding $\langle\hat{\mathbf{S}}_{i}^{\text{sp-el}}\rangle$. Yet, since $\eta^{\prime}=0.14$, the sp-el and sp ground states noticeably differ. Altogether, these comparisons confirm that i-electrons markedly affect the quantum nanoskyrmion and that it is beneficial, even with a mean-field treatment of Kondo-like exchange, to have quantum rather than classical l-spins.
@@
Skyrmion ground states. To simulate nanosized skyrmions in isolated/open systems we take $D/J\sim 2-5$. These are large values, but consistent with the giant DMI ones from some monolayer materials [65]. The other parameters used, $J/t_{h}\sim 0.1-0.5$ and $g/t_{h}\sim 1-2$, are also congruous with typical values in the literature [45].
@@
Isolated system.- We consider is a $21\times 21$ rhombus cluster (which permits to simulate a skyrmion lattice) with OBC and parameters $D=0.2,J=0.1,B^{z}=0.1,t_{h}=1,g=2$. The ground state profile of such spin+electron system is shown in Fig. 3a, where the skyrmions exhibit an approximate periodic alignment, whilst the sites with $S^{z}\simeq\frac{1}{2}$ state (black circles) can be thought of as domain walls. Since the l-spins are described quantum mechanically, we can look at entanglement between spin pairs at sites $i\neq j$ [66, 67, 48], quantified by the concurrence $\mathcal{C}_{ij}=\mathcal{C}[\hat{\rho}_{ij}]$, and where $\hat{\rho}_{ij}$ is the reduced density matrix. Similarly to what found in [48], $\mathcal{C}$ shows a lack of long-range entanglement for the skyrmion (crystal) texture. However, since the distance $\eta^{\prime}$ between sp-el and sp systems is $10\%$, the concurrences with and without i-electrons discernibly differ. In turn, the i-electrons are affected by the skyrmion texture (for example, the local density of states for spin-up and spin-down i-electrons show large (small) imbalance at (far away from) the skyrmions-core sites, see the SM).
@@
Further insight comes from the spin structure factor, defined as $S^{\alpha\beta}(\mathbf{k})=N^{-1}\sum_{ij\in C}\langle\hat{S}_{i}^{\alpha}\hat{S}_{j}^{\beta}\rangle e^{i\mathbf{k}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}$, and with $\alpha\beta$ denoting either the $zz$ or the ++-+ - components. In Fig. 3b (for clarity, we show $\ln S^{\alpha\beta}$ rather than $S^{\alpha\beta}$), both $S^{zz}({\bf k})$ and $S^{+-}({\bf k})$ exhibit sixfold intensity patterns, consistent with neutron scattering results (from e.g. the B20 compound MnSi [3]). Overall, the signal is stronger for $S^{zz}$ than for $S^{+-}$; in particular, $S^{zz}$ has a bright spot at ${\bf k}=0$, due to the strong degree of spin orientation along $z$. Additionally, the logarithmic display reveals additional but much less intense, lower symmetry features, that we ascribe to size and the open-boundary-conditions effects in the rhombus cluster.
@@
Ground state for an open system. Using NEGF for the i-electrons and Tensor networks for the l-spins, we have also considered the case of a $6\times 6$ sp-el square cluster attached to leads. The isolated central region ground state at $t=0$ (determined in the mean-field Kondo-like exchange) is connected to the leads adiabatically for $500<t<$ 50. The evolution of the sp-el system can then be continued to $t_{0}=125$, which establishes steady spin-polarised currents through the central region in equilibrium. These currents are weak and essentially present only at the boundary sites of $C$ (see the SM). The ground state average spin distributions for this setup, with $D=1,J=0.2,B^{z}=0.5,t_{h}=1,g=1$, are shown in Fig. 3c. They result in a value $\eta=0.06$, i.e. a non-strong average influence of the electrons on the spin-texture. Still, an appreciable and interesting feature is that the spins of the i-electrons are tilted compared to the isolated case (not shown). Since for the open case there is a spin-polarized electric current flowing (and a corresponding induced magnetic field), we attribute the spin tilt to an induced effective spin-orbit interaction.
@@
Dynamics. For out-of-equilibrium situations, the natural option would be to consider the open system of Fig. 1iii where, after reaching equilibrium at $t=t_{0}$ (as discussed above, via a slow-ramping dynamics to connect the leads and dissipate fluctuations), time-dependent currents are injected in the electron subsystem via a bias. For example, for spin-up polarization and with reference to Eq. (2), $\hat{H}^{s}\rightarrow\hat{H}^{s}(t)\equiv\hat{H}^{s}+\hat{w}_{L}(t)+\hat{w}_{R}(t)$, where $\hat{w}_{L}(t)=\sum_{i\in L}\epsilon_{L}(t)\hat{c}^{\dagger}_{i\uparrow}\hat{c}_{i\uparrow}$ and similar for $\hat{w}_{R}(t)$. The time evolution would then be performed via NEGF for the electrons and via e.g. TEBD for the spins.
@@
We defer the study of the NEGF+TEBD dynamics of a l-spins+i electrons system connected to semi-infinite leads to a forthcoming paper. Here, as proof of concept of our approach, we use the enlarged isolated setup of Fig. 2b, with a region $C$ consisting of a $4\times 10$ rhombus cluster, and an enlarged electron region of $4\times(100+10+100)$ sites, to delay the reflection of the currents by the boundaries, and to have steady, stable currents established within the time of interest. The system is considered within OBC. In our simulations, we set $t_{0}\!=\!0,D\!=\!0.2,J\!=\!0.05,B^{z}\!=\!0.06,t_{h}\!=\!1,g\!=\!2$. In this case, the QQmf ground state $|\Psi^{\text{el}}(0)\rangle$ (obtained without time propagation, and shown at $t\!=\!0$ in Fig. 4b) corresponds to a spin texture with multiple meron-like (i.e. half-skyrmion-like) structures. These (henceforth called merons for simplicity) are centered at sites $i$ of the cluster boundaries where $\langle\hat{S}_{i}^{z}\rangle<0$, with $\langle\hat{S}_{i^{\prime}}^{z}\rangle>0$ for their second and third neighbours sites $i^{\prime}$. To support this interpretation and address the individual behavior of merons, we use a modified, “local” scalar chirality $Q_{i}$ calculated over all the triangles including sites $i$ and $i^{\prime}$ (for example, at $t\!=\!0$, $Q_{i}\sim-0.16$ for the green shaded meron in Fig. 4b, see the SM for details).
@@
Then, starting from $|\Psi^{\text{el}}(0)\rangle$, we introduce for $t>0$ a bias $\epsilon_{L}(t)=0.5[1-\theta(10-t)\cos^{2}(\pi t/20)]$ with $\epsilon_{R}(t)=-\epsilon_{L}(t)$. At each time step of the time evolution, $\hat{H}^{d}_{MF}$ is constructed with input from the time-evolved spin expectation $\langle\hat{\mathbf{s}}_{i}\rangle_{t}$, and the l-spins are evolved via TEBD. Subsequently, an updated $\langle\hat{\mathbf{S}}_{i}\rangle_{t}$ and the bias $\epsilon_{L/R}(t)$ drive the electrons via $\hat{H}^{s}_{MF}(t)$, and the full spin-electron system is thus evolved for one time step.
@@
Figure 5: Wmf (top) vs QQmf (bottom) structure factor $S^{\alpha\beta}(\mathbf{k})$ at $t=125$ for the $4\times(100+10+100)$ $i$-electron+$l$-spin texture. Model parameters are $D=0.2,J=0.05,B^{z}=0.06,t_{h}=1,g=2$. Spin-up polarized bias $\epsilon_{L}=-\epsilon_{R}$ in the enlarged region switched-on at $t\!=\!0$ and ramped to maximum strength 0.5 at $\!t=\!10$.
@@
The time evolution of selected spin-up bond currents $I_{\langle ij\rangle}=-2t_{h}\text{Im}\langle\hat{c}_{j\uparrow}^{\dagger}\hat{c_{i\uparrow}}\rangle$, is shown in Fig. 4a. After the transient regime ($t\lesssim 10$), the currents at/near the edges ($I_{AF,AE,AB}$) stay rather stable till $t\approx 80$; then, they start to be reflected at the outer boundaries of the $L,R$ regions. By contrast, currents away from the edges ($I_{BE,BD,BG}$) undergo reflection sooner (i.e., $t\approx 50$). In Fig. 4a, the currents at different bonds differ for magnitude and/or sign (also for parallel adjacent bonds). This is because of the merons: these are driven in a direction opposite (see the shaded area in the time snapshots) to the average flow of the spin-up electrons (also, small size and nonequilibrium quantum fluctuations make the merons to slightly change their shape in time). The i-electrons, in turn, conform their flow to the presence of the moving merons, which results in the observed behavior of the bond spin-up currents. Furthermore, in the ground state ($t=0$), the i-electron’s spins texture mirrors the l-spins one, but in a much weaker pattern. The latter extends outside region $C$, while fading as the distance from $C$ increases. Yet, with currents (i.e. $t>0$), the outer parts of the pattern move towards $C$, affecting the l-spins texture. This is why, for example, $\langle\hat{S}^{z}\rangle$ at the top-left site in Fig. 4b changes sign during the time evolution. To further qualify the QQmf approach and the inherent results, it can be useful to employ the Wmf approach for the present system and situation. In doing so, we find that the local spin expectation values from the QQmf and Wmf methods are initially quite similar ($\eta(t=0)=3\%$), but their difference increases during time evolution ($\eta(t=125)=7\%)$. We also examined the spin-spin correlation functions $S^{\alpha\beta}(\mathbf{k})$ (within Wmf, $\langle\hat{S}^{\alpha}_{i}\hat{S}^{\beta}_{j}\rangle\approx\langle\hat{S}^{\alpha}_{i}\rangle\langle\hat{S}^{\beta}_{j}\rangle$). These look different in the two descriptions, as illustrated for $S^{\alpha\beta}(\mathbf{k}(t=125))$ in Fig. 5 (and for $S^{\alpha\beta}(\mathbf{k}(t=0))$ in the SM). These trends align with the typical behavior of mean-field treatments in condensed matter: while corrections to mean-field values may often be small for some observables, they can still produce noticeable differences in some physical quantities.
@@
Conclusions. We have proposed a computationally viable theoretical approach to describe quantum nanoskyrmions for in- and out-of-equilibrium itinerant-electron+localised-spin systems, with scope and results different from classical-spin+quantum-electron or spin-only treatments. It combines Matrix Product States methods for the localised-spins, with exact diagonalization or nonequilibrium Green’s function methods for the itinerant electrons, while treating the spin-electron interaction at the mean-field level.
@@
Besides nanoskyrmions, the method permits to address a vast range of physical phenomena in (topological or not, ordered or not) magnetic systems that are elusive to classical-spin or spin-only schemes, for example magnetoresistance due to quantum spin-textures, the photoexcitation of quantum skyrmions, or the photoelectron spectrum from the itinerant electrons in the presence of quantum spin textures). Furthermore, the scheme is amenable to several extensions, and thus likely expected to foster new avenues of research. As next developments, we are considering ways to go beyond a mean-field treatment of the spin-electron interaction via non perturbative local approximations, and the inclusion photon-skyrmion interactions via Tensor Networks [68, 69] and nonequilibrium Green’s function formulations [70, 35, 71].

### v4

@@ Abstract
Next, to discuss the role of the i electrons, we move to a spin-electron (sp-el) and a spin-only (sp) system in a rhombus-shaped $11\times 1$ 11 finite triangular lattice, also treated with OBC. We again compare their respective ground states using $\eta^{\prime}=N^{-1}\sum_{i\in C}\big{|}\langle\hat{\mathbf{S}}_{i}^{\text{sp-el}}\rangle-\langle\hat{\mathbf{S}}_{i}^{\text{sp}}\rangle\big{|}\big{/}\big{|}\langle\hat{\mathbf{S}}_{i}^{\text{sp}}\rangle\big{|}$. We choose $D\!\!=\!\!0.2,J\!\!=\!\!0.15,B^{z}\!\!=\!\!0.06,t_{h}\!\!=\!\!1,g\!\!=\!\!2$, for which both systems exhibit a single-skyrmion texture. This is confirmed by the value of the quantum scalar chirality [44] $Q=\pi^{-1}\sum_{<ijk>\in C}\langle\hat{\mathbf{S}}_{i}\cdot[\hat{\mathbf{S}}_{j}\times\hat{\mathbf{S}}_{k}]\rangle$, where the sum runs over all nonoverlapping triangles formed by neighboring sites $i$, $j$, $k$. For $S=1/2$ l spins, $|Q|\sim 1$ signals the presence of a skyrmion. For the case at hand we respectively find $Q_{\text{sp}}^{\text{sp-el}}=-0.618$, and $Q^{\text{sp}}=-0.624$. Even in this strong Kondo regime ($g/t_{h}=2$), the mean-field electron-spin expectation values $\langle\hat{\mathbf{s}}_{i}\rangle$ are only a few percent of the corresponding $\langle\hat{\mathbf{S}}_{i}^{\text{sp-el}}\rangle$. Yet, since $\eta^{\prime}=0.14$, the sp-el and sp ground states noticeably differ. Altogether, these comparisons confirm that i electrons markedly affect the quantum nanoskyrmion and that it is beneficial, even with a mean-field treatment of Kondo-like exchange, to have quantum rather than classical l spins.
@@
Skyrmion ground states. To simulate nanosized skyrmions in isolated or open systems we take $D/J\sim 2-5$. These are large values, but consistent with the giant DMI ones from some monolayer materials [66]. The other parameters used, $J/t_{h}\sim 0.1-0.5$ and $g/t_{h}\sim 1-2$, are also congruous with typical values in the literature [41].
@@
Isolated system. We consider a $21\times 21$ rhombus cluster (which permits to simulate a skyrmion lattice) with OBC and parameters $D=0.2,J=0.1,B^{z}=0.1,t_{h}=1,g=2$. The ground state profile of such spin+electron system is shown in Fig. 3(a), where the skyrmions exhibit an approximate periodic alignment, while the sites with $S^{z}\simeq\frac{1}{2}$ state (black circles) can be thought of as domain walls. Since the l spins are described quantum mechanically, we can look at entanglement between spin pairs at sites $i\neq j$ [67, 68, 48], quantified by the concurrence $\mathcal{C}_{ij}=\mathcal{C}[\hat{\rho}_{ij}]$, and where $\hat{\rho}_{ij}$ is the reduced density matrix. Similarly to what is found in [48], $\mathcal{C}$ shows a lack of long-range entanglement for the skyrmion (crystal) texture. However, since the distance $\eta^{\prime}$ between sp-el and sp systems is $10\%$, the concurrences with and without i electrons discernibly differ. In turn, the i electrons are affected by the skyrmion texture (for example, the local density of states for spin-up and spin-down i electrons show large (small) imbalance at (far away from) the skyrmions-core sites, see the SM).
@@
Further insight comes from the spin structure factor, defined as $S^{\alpha\beta}(\mathbf{k})=N^{-1}\sum_{ij\in C}\langle\hat{S}_{i}^{\alpha}\hat{S}_{j}^{\beta}\rangle e^{i\mathbf{k}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}$, and with $\alpha\beta$ denoting either the $zz$ or the ++-+ - components. In Fig. 3(b) (for clarity, we show $\ln S^{\alpha\beta}$ rather than $S^{\alpha\beta}$), both $S^{zz}({\bf k})$ and $S^{+-}({\bf k})$ exhibit sixfold intensity patterns, consistent with neutron scattering results (from e.g. the B20 compound MnSi [3]). Overall, the signal is stronger for $S^{zz}$ than for $S^{+-}$; in particular, $S^{zz}$ has a bright spot at ${\bf k}=0$, due to the strong degree of spin orientation along $z$. Additionally, the logarithmic display reveals additional but much less intense, lower symmetry features, that we ascribe to size and the open-boundary-conditions effects in the rhombus cluster.
@@
Ground state for an open system. Using NEGF for the i electrons and tensor networks for the l spins, we have also considered the case of a $6\times 6$ sp-el square cluster attached to leads. The isolated central region ground state at $t=0$ (determined in the mean-field Kondo-like exchange) is connected to the leads adiabatically for $500<t<$ 50. The evolution of the sp-el system can then be continued to $t_{0}=125$, which establishes steady spin-polarized currents through the central region in equilibrium. These currents are weak and essentially present only at the boundary sites of $C$ (see the SM). The ground state average spin distributions for this setup, with $D=1,J=0.2,B^{z}=0.5,t_{h}=1,g=1$, are shown in Fig. 3(c). They result in a value $\eta=0.06$, i.e., a nonstrong average influence of the electrons on the spin texture. Still, an appreciable and interesting feature is that the spins of the i-electrons are tilted compared to the isolated case (not shown). Since for the open case there is a spin-polarized electric current flowing (and a corresponding induced magnetic field), we attribute the spin tilt to an induced effective spin-orbit interaction.
@@
Dynamics. For out-of-equilibrium situations, the natural option would be to consider the open system of Fig. 1iii where, after reaching equilibrium at $t=t_{0}$ (as discussed above, via a slow-ramping dynamics to connect the leads and dissipate fluctuations), time-dependent currents are injected in the electron subsystem via a bias. For example, for spin-up polarization and with reference to Eq. (2), $\hat{H}^{s}\rightarrow\hat{H}^{s}(t)\equiv\hat{H}^{s}+\hat{w}_{L}(t)+\hat{w}_{R}(t)$, where $\hat{w}_{L}(t)=\sum_{i\in L}\epsilon_{L}(t)\hat{c}^{\dagger}_{i\uparrow}\hat{c}_{i\uparrow}$ and similar for $\hat{w}_{R}(t)$. The time evolution would then be performed via NEGF for the electrons and via, e.g., TEBD for the spins.
@@
We defer the study of the NEGF+TEBD dynamics of a l spins+i electrons system connected to semi-infinite leads to a forthcoming paper. Here, as proof of concept of our approach, we use the enlarged isolated setup of Fig. 2(b), with a region $C$ consisting of a $4\times 10$ rhombus cluster, and an enlarged electron region of $4\times(100+10+100)$ sites, to delay the reflection of the currents by the boundaries, and to have steady, stable currents established within the time of interest. The system is considered within OBC. In our simulations, we set $t_{0}\!=\!0,D\!=\!0.2,J\!=\!0.05,B^{z}\!=\!0.06,t_{h}\!=\!1,g\!=\!2$. In this case, the QQmf ground state $|\Psi^{\text{el}}(0)\rangle$ [obtained without time propagation, and shown at $t\!=\!0$ in Fig. 4(b)] corresponds to a spin texture with multiple meron-like (i.e. half-skyrmion-like) structures. These (henceforth called merons for simplicity) are centered at sites $i$ of the cluster boundaries where $\langle\hat{S}_{i}^{z}\rangle<0$, with $\langle\hat{S}_{i^{\prime}}^{z}\rangle>0$ for their second and third neighbors’ sites $i^{\prime}$. To support this interpretation and address the individual behavior of merons, we use a modified, “local” scalar chirality $Q_{i}$ calculated over all the triangles including sites $i$ and $i^{\prime}$ (for example, at $t\!=\!0$, $Q_{i}\sim-0.16$ for the green shaded meron in Fig. 4(b); see the SM for details).
@@
Then, starting from $|\Psi^{\text{el}}(0)\rangle$, we introduce for $t>0$ a bias $\epsilon_{L}(t)=0.5[1-\theta(10-t)\cos^{2}(\pi t/20)]$ with $\epsilon_{R}(t)=-\epsilon_{L}(t)$. At each time step of the time evolution, $\hat{H}^{d}_{\text{MF}}$ is constructed with input from the time-evolved spin expectation $\langle\hat{\mathbf{s}}_{i}\rangle_{t}$, and the l spins are evolved via TEBD. Subsequently, an updated $\langle\hat{\mathbf{S}}_{i}\rangle_{t}$ and the bias $\epsilon_{L/R}(t)$ drive the electrons via $\hat{H}^{s}_{\text{MF}}(t)$, and the full spin-electron system is thus evolved for one time step.
@@
Figure 5: Wmf (top) vs QQmf (bottom) structure factor $S^{\alpha\beta}(\mathbf{k})$ at $t=125$ for the $4\times(100+10+100)$ $i$ electron+$l$ spin texture. Model parameters are $D=0.2,J=0.05,B^{z}=0.06,t_{h}=1,g=2$. Spin-up polarized bias $\epsilon_{L}=-\epsilon_{R}$ in the enlarged region switched on at $t\!=\!0$ and ramped to maximum strength 0.5 at $\!t=\!10$.
@@
The time evolution of selected spin-up bond currents $I_{\langle ij\rangle}=-2t_{h}\text{Im}\langle\hat{c}_{j\uparrow}^{\dagger}\hat{c_{i\uparrow}}\rangle$ is shown in Fig. 4(a). After the transient regime ($t\lesssim 10$), the currents at or near the edges ($I_{AF,AE,AB}$) stay rather stable until $t\approx 80$; then, they start to be reflected at the outer boundaries of the $L,R$ regions. By contrast, currents away from the edges ($I_{BE,BD,BG}$) undergo reflection sooner (i.e., $t\approx 50$). In Fig. 4(a), the currents at different bonds differ for magnitude and/or sign (also for parallel adjacent bonds). This is because of the merons: these are driven in a direction opposite (see the shaded area in the time snapshots) to the average flow of the spin-up electrons (also, small size and nonequilibrium quantum fluctuations make the merons slightly change their shape in time). The i electrons, in turn, conform their flow to the presence of the moving merons, which results in the observed behavior of the bond spin-up currents. Furthermore, in the ground state ($t=0$), the i electron’s spins texture mirrors the l spins one, but in a much weaker pattern. The latter extends outside region $C$, while fading as the distance from $C$ increases. Yet, with currents (i.e. $t>0$), the outer parts of the pattern move towards $C$, affecting the l spins texture. This is why, for example, $\langle\hat{S}^{z}\rangle$ at the top-left site in Fig. 4(b) changes sign during the time evolution. To further qualify the QQmf approach and the inherent results, it can be useful to employ the Wmf approach for the present system and situation. In doing so, we find that the local spin expectation values from the QQmf and Wmf methods are initially quite similar [$\eta(t=0)=3$], but their difference increases during time evolution [$\eta(t=125)=7$]. We also examined the spin-spin correlation functions $S^{\alpha\beta}(\mathbf{k})$ (within Wmf, $\langle\hat{S}^{\alpha}_{i}\hat{S}^{\beta}_{j}\rangle\approx\langle\hat{S}^{\alpha}_{i}\rangle\langle\hat{S}^{\beta}_{j}\rangle$). These look different in the two descriptions, as illustrated for $S^{\alpha\beta}[\mathbf{k}(t=125)]$ in Fig. 5 (and for $S^{\alpha\beta}[\mathbf{k}(t=0)]$ in the SM). These trends align with the typical behavior of mean-field treatments in condensed matter: while corrections to mean-field values may often be small for some observables, they can still produce noticeable differences in some physical quantities.
@@
Conclusions. We have proposed a computationally viable theoretical approach to describe quantum nanoskyrmions for in- and out-of-equilibrium itinerant-electron+localized-spin systems, with scope and results different from classical-spin+quantum-electron or spin-only treatments. It combines matrix product states methods for the localized-spins, with exact diagonalization or nonequilibrium Green’s function methods for the itinerant electrons, while treating the spin-electron interaction at the mean-field level.
@@
Besides nanoskyrmions, the method permits one to address a vast range of physical phenomena in (topological or not, ordered or not) magnetic systems that are elusive to classical-spin or spin-only schemes, for example, magnetoresistance due to quantum spin textures, the photoexcitation of quantum skyrmions, or the photoelectron spectrum from the itinerant electrons in the presence of quantum spin textures. Furthermore, the scheme is amenable to several extensions, and thus likely expected to foster new avenues of research. As next developments, we are considering ways to go beyond a mean-field treatment of the spin-electron interaction via nonperturbative local approximations, and the inclusion photon-skyrmion interactions via tensor networks [69, 70] and nonequilibrium Green’s function formulations [71, 35, 72].